OpenStudy (anonymous):

what is the sum of the first 4 terms of the arithmetic sequence in which the 6th term is 8 and the 10th term is 13?

OpenStudy (campbell_st):

the general equation of a term in an arithmetic sequence is \[T_{n} = a + (n - 1)d\] so setting up the equations you will have \[8 = a + (6 -1) d\] and \[13 = a + (10 -1) d\] simplifying these equations you have \[8 = a + 5d\] and \[13 = a + 9d\] solve simultaneously using substitution let a = 8 - 5d then 13 = (8 - 5d) + 9d or 13 = 8 + 4d d = 1.25 substitute to find a a = 1.75 now you have a, d and you know n = 4... so the general equation of the sum of an arithmetic sequence is \[S_{n} = \frac{n}{2}[2a + (n -1)d]\] just substitute and solve.